# Bridging the Culture Gap: A Framework for LLM-Driven Socio-Cultural Localization of Math Word Problems in Low-Resource Languages

### v2

@@ 1 Introduction
Low-resource language (LRL) math word problem evaluation heavily relies on human or machine-translated benchmarks, as demonstrated by the translation of GSM8K Cobbe et al. (2021) (a dataset of 8.79K high-quality, linguistically diverse grade school math word problems) into GSM8K_zh for Chinese Yu et al. (2023), MGSM for ten typologically diverse languages Shi et al. (2022a), and AfriMGSM for 18 African languages Adelani et al. (2024). Existing human- or machine-translated benchmarks often fail to account for local cultural contexts that contain day to day activities, common names and currencies. As a result, evaluating LLMs on such non-localized data primarily measures their mathematical understanding ability in English centric scenarios, names and currencies.
@@
  • We develop an LLM-driven localization pipeline to generate culturally adapted versions of translated datasets by replacing key entities with apt socio-cultural variants. [^1] https://huggingface.co/collections/israel/cultural-localilzation-687e392f6be2aaa0b7c7831c
@@ 5 Entity replaced
$x_{\text{ent}}=fn(x_{\text{ent}},replacement\_dict)$ Camari owes Julani shilingi 100. They agreed to have monthly interest of 2%. If Camari was able to pay it after 3 months, how much should she give to Julani?
@@ 2 Related Work
Math word problems are mathematical questions framed in everyday language, often grounded in real-life scenarios and activities rather than expressed purely through abstract symbols or equations Cobbe et al. (2021). These problems serve as an engaging way to learn mathematical reasoning, as they require not only computational skills but also the ability to interpret and model real-world situations.
LLM Robustness.
@@
Although LLMs are highly capable of solving complex tasks, they often struggle with simple variations in input (Askari et al., 2025; Chhabra et al., 2024; Achara and Chhabra, 2025). For example, Abedin et al. (2025) demonstrate that even minor perturbations such as spelling mistakes can significantly impact model performance, a challenge that is particularly pronounced in low-resource settings and evaluations.
Translation and Localization.
@@
Translating existing benchmarks is one of the simplest approaches to benchmark creation, particularly for low-resource languages that lack dedicated evaluation datasets. In such cases, researchers often resort to translating established benchmarks from high-resource languages to enable evaluation and comparison Adelani et al. (2024); Koto et al. (2024); Li et al. (2023); Son et al. (2024). In African context close to 30% of resource papers translate existing benchmarks to create language specific benchmarks Alabi et al. (2025). However, to better reflect native cultural identities and move away from the Western-centric concepts that often persist in translated benchmarks, some researchers have begun developing fully localized benchmarks Yu et al. (2025). This work complements manual localization and cultural adaptation efforts by reducing the scale and resources required to acquire such data. Just as traditional data augmentation methods improve model robustness without the need for entirely new data creation, our proposed approach similarly complements translation by enabling the creation of large-scale, culturally adapted datasets with minimal overhead.
Automatic Entity based Augmentation.
@@
In addition to manual localization efforts carried out by volunteers, Ye et al. (2024) introduce a novel data augmentation technique that leverages large language models (LLMs). Specifically, they apply this method to enhance performance in few-shot Named Entity Recognition (NER) tasks, demonstrating that LLM-driven augmentations can serve as a valuable complement to human-curated resources.
Mathematics and Cultural Entities.
@@
Karim et al. (2025); Tomar et al. (2025) examined the influence of cultural context on mathematical problems by analyzing the impact of culturally grounded entities such as personal names and food items. However, one critical dimension that remains underexplored is the role of language itself and the biases that can be propagated through translation. Since the significant number of multilingual training and evaluation datasets are created via translation from high-resource languages Alabi et al. (2025), they often fail to capture the linguistic and cultural nuances of the target languages. Evaluating model robustness in the presence of culturally specific entities is a critical challenge, especially as LLMs are increasingly deployed in real-world applications. Native language speakers naturally refer to familiar names, organizations, and currencies from their own cultural context, making it essential for models to handle such variations reliably.
@@ 3.1 Pipeline
As shown in stage 7 of Table 1, our quality control measures ensure no overlapping or inconsistent entity replacements, enforce a single currency type per problem to avoid conversion errors, verify that localized outputs match the length of their direct translations, and check similarity between localized and translated text using the difflib library333docs.python.org/3/library/difflib.html. Since all languages share the same entity-replaced text ($x_{\text{ent}}$), human verification only requires comparing entities in the replacement dictionary. This prevents unnecessary additions by LLMs, maintains prompt consistency, and keeps outputs close to the original structure.
@@ 3.2 Datasets Used
Evaluation Dataset. AfriMGSM Adelani et al. (2024) is a manually translated benchmark spanning 18 languages sourced from MGSM Shi et al. (2022b). We found that approximately 86% of the test set, includes at least one important entity. Using the English and manually translated pairs, we created a localized evaluation dataset by replacing these entities with culturally appropriate local alternatives while keeping the rest of the problem content unchanged. The correctness of the localized dataset was verified through manual inspection by the authors. Since the approach returns the unlocalized item if it fails, no additional noise is introduced.
@@
To ensure translation quality, we filtered the outputs using SSA-COMMET Li et al. (2025), a sentence-level semantic similarity metric that scores translation quality on a scale from 0 to 1, with 1 indicating perfect translation. Based on the score distribution shown in the Appendix D, we retained only translations with a COMMET score above 0.65. From these, we selected the top 1,500 examples per language for automatic localization and use as training data in our after localizing each. While intermediate experiments leverage Gemini-1.5-pro to evaluate the pipeline and assess its effectiveness due to cost constraints and to avoid unwanted behaviors, such as prompt caching [^4] https://ai.google.dev/gemini-api/docs/caching?lang=python, the final dataset is generated using Gemini-2.5-pro to ensure the highest quality standards.
Figure 3: Direct Translation vs. Auto Localization performance for Aya-expanse-32b (top left), LLaMA-3-70B-Instruct (top right), Gemma-2-27b-it (bottom left), and GPT-4o-mini (bottom right). We observe that some languages, such as Swahili, are more prone to confusion from newly introduced entities in three out of four models, highlighting the lack of robustness for local entity presence in low-resource languages.
Figure 4: Impact of Cultural Entities on English benchmarks. Does the presence of cultural entities ($x_{\text{ent}}$; see Table 1) affect model performance? Results are shown for Aya-expanse-32b (top-left), LLaMA-3-70B-Instruct (top-right), Gemma-2-27b-it (bottom-left), and GPT-4o-mini (bottom-right). We observe that the presence of local entities has a consistent effect on evaluations, with the exception of Gemma-2-27b-it.
@@ 4 Experimental Setup
Evaluated LLMs. In this work, we evaluated both open-source and closed-source models commonly studied in existing multilingual mathematical research. Our selection spans a range of model sizes, including Aya-expanse-32b Dang et al. (2024), LLaMA-3-8B-Instruct, LLaMA-3-70B-Instruct Grattafiori et al. (2024), Gemma-2-9b-it, Gemma-2-27b-it Team et al. (2024) and GPT-4o-mini. The finetuning experiments are done only on LLaMA-3-8B-Instruct, Gemma-2-9b-it because of compute constraints.
Evaluation Metrics. Most existing work on mathematical word problems focuses on evaluating reasoning capabilities, where intermediate calculation steps are used as part of the input or as supervision for general benchmarking. However, due to the lack of multilingual reasoning benchmarks and the challenges involved in extracting step-by-step reasoning across 18 different languages, this research focuses solely on evaluating final answers.
Several metrics are commonly used in mathematical evaluation research, including exact match, numeric match. In this work, we select the metrics that best align with the core objectives of our study. Exact Match (EM) evaluates whether the predicted answer exactly matches the reference answer as a string. Numeric Match (NM) checks whether the predicted numerical value matches the ground truth, ignoring differences in formatting such as units or punctuation after converting them into floating point data and account for errors between them. Moreover, for ensuring robustness and minimal noise via prompt selection, we adopted three prompt variations from Adelani et al. (2024) for our experiments. We further customized them to return only the final output without intermediate steps and reported the average score across all three prompts.
Reproducibility. We used ElutherAI’s open source Language Model Evaluation Harness (lm-eval) framework Gao et al. (2024) to evaluate models. Instead of direct model querying this allows us to have standardized reproducible evaluations across all LLMs. We also set generation parameters (i.e. temperature) to zero for consistency.
Language Translated English with Entities Auto Localized All Data
$x_{trans}$ $x_{ent}$ $\hat{x}_{\text{loc}}$
[…]
# Native Robust lang. 6 4 7 11
Language Translated English with Entities Auto Localized All Data
$x_{trans}$ $x_{ent}$ $\hat{x}_{\text{loc}}$
[…]
# Native Robust lang. 7 8 12 10
Figure 5: Native Robustness for Model (Numeric Match $\Delta$). For each language, we report $\Delta_{\mathrm{NM}}=\mathrm{NM}_{\text{localized}}-\mathrm{NM}_{\text{translated}}$ after LoRA fine-tuning of LLaMA-3-8B-Instruct (left) and Gemma-2-9b-it (right) across training-data configurations. Positive denotes higher accuracy on localized benchmarks (greater robustness to local entities); Negative denotes higher accuracy on English-centric benchmarks (reduced robustness to localization) and Yellow denotes that the model performance did not degrade or improve. Languages are ordered from lower- to higher-resource by FineWeb2 coverage Penedo et al. (2025).
@@ 5 Results and Analysis
(RQ1) Can LLM-driven pipelines be used to localize translated benchmarks?
To address the high cost associated with using human translators for creating localized versions of datasets, we introduce an automatic pipeline that generates culturally adapted versions of mathematical word problems. While this pipeline is not intended to replace human annotators, it serves as a valuable tool for producing augmented datasets that help improve the robustness of LLMs to entity-based variations in problems.
In our first set of experiments (please see Figure 2), we used Gemini-1.5-pro to generate localized versions of the dataset using both our six-stage localization pipeline and direct prompting via the Gemini API. Once the two versions of the localized dataset were generated, three annotators independently labeled each instance as either a valid or invalid localization. Additionally we wanted to show this method does not require the latest/largest models – allowing us to achieve stellar performance even with older models like Gemini-1.5-pro.
@@
(RQ2) Does introducing local entities in translated benchmarks reveal performance disparities for MWPs in LRLs and English?
We first analyze LRLs and then English language MWPs. Given that 30% of newly created African datasets are based on translated content Alabi et al. (2025), it is important to assess whether LLMs are overfitting to English-centric entities commonly preserved during translation. Such overfitting may lead to performance degradation when these entities are replaced with their native counterparts. In this work, we hypothesize that models should perform better when evaluated on data containing entities from their target (native) language, as opposed to data where entities originate from other languages.
Figure 3 illustrates the performance differences between models evaluated on translated benchmarks and those evaluated on automatically localized benchmarks, highlighting the impact of entity localization on model robustness. We can observe that the Aya-expanse-32b LLM performs well on translated benchmarks while LLaMA-3-70B-Instruct attains a performance gain from seeing native entities. Gemma-2-27b-it has more consistent results across native and English centric entities.
@@
Both LLaMA-3-70B-Instruct and GPT-4o-mini achieve higher accuracy on the localized (native-entity) variant than on the English benchmark, whereas the Aya-expanse-32b model’s performance lags behind. Gemma-2-27b-it generally tracks the localized variant more closely, despite occasional dips and overshoots.
@@
Figure 5 presents the performance of models trained on different flavors of MWP data. We begin with models trained on the translated dataset, followed by those trained on the original English GSM8K dataset, and then models trained using the training data produced by our localization pipeline. Finally, we include results for models trained on a combined dataset that includes all three sources.
Our evaluation compares model performance on the localized benchmark versus the translated benchmark. This comparison helps us understand whether the inclusion of native cultural entities affects model behavior. A positive difference indicates that the model performs better on the localized benchmark, suggesting improved robustness to cultural variations. Conversely, a negative difference highlights cases where the model handles English-specific entities such as names and currencies effectively but fails when presented with culture-specific alternatives.
We observe that both LLaMA-3-8B-Instruct and Gemma-2-9b-it models exhibit improved robustness to entity changes in several languages. For LLaMA-3-8B-Instruct, the best performance is achieved when localization is combined with additional noisy data, indicating that diverse training sources can enhance generalization. In contrast, for Gemma-2-9b-it, models trained exclusively on Auto Localization data perform best, suggesting that this model benefits more from clean, culturally adapted input.
Additionally, we find that languages with more training resources tend to benefit more from localization. This trend is evident from the table, where languages are sorted in ascending order based on their availability in the FineWeb2 Penedo et al. (2025) dataset. Higher-resource languages (appearing lower in the table) show greater performance gains, while lower-resource languages (toward the top) experience more limited improvements. More exploration to see characteristics of languages affected by the finetuning experiment included in the Appendix C.

### v3

@@ 1 Introduction
Low-resource language (LRL) math word problem evaluation heavily relies on human or machine-translated benchmarks, as demonstrated by the translation of GSM8K Cobbe et al. (2021) (a dataset of 8.79K high-quality, linguistically diverse grade school math word problems) into GSM8K_zh for Chinese Yu et al. (2023), MGSM for 10 typologically diverse languages Shi et al. (2022a), and AfriMGSM for 18 African languages Adelani et al. (2024). Existing human- or machine-translated benchmarks often fail to account for local cultural contexts, which include day-to-day activities, common names, and local currencies. As a result, evaluating LLMs on such non-localized data primarily measures their mathematical understanding ability in English-centric scenarios, names, and currencies.
@@
  • We develop an LLM-driven localization pipeline to generate culturally adapted versions of translated datasets by replacing key entities with apt socio-cultural variants and processed localized MWP datasets. [^1] https://github.com/IsraelAbebe/Auto-Localizer.
@@ 5 Entity replaced
$x_{\text{ent}}=fn(x_{\text{trans}},replacement\_dict)$ Camari owes Julani shilingi 100. They agreed to have monthly interest of 2%. If Camari was able to pay it after 3 months, how much should she give to Julani?
@@ 2 Related Work
Math World Problems. Math word problems are mathematical questions framed in everyday language, often grounded in real-life scenarios and activities rather than expressed purely through abstract symbols or equations Cobbe et al. (2021). These problems serve as an engaging way to learn mathematical reasoning, as they require not only computational skills but also the ability to interpret and model real-world situations.
@@
LLM Robustness. Although LLMs are highly capable of solving complex tasks, they often struggle with simple variations in input (askari2025assessing; chhabra2024revisiting; achara2025watching). For example, Abedin et al. (2025) demonstrate that even minor perturbations such as spelling mistakes can significantly impact model performance, a challenge that is particularly pronounced in low-resource settings and evaluations.
@@
Translation and Localization. Translating existing benchmarks is one of the simplest approaches to benchmark creation, particularly for low-resource languages that lack dedicated evaluation datasets. In such cases, researchers often resort to translating established benchmarks from high-resource languages to enable evaluation and comparison Adelani et al. (2024); Koto et al. (2024); Li et al. (2023); Son et al. (2024). In African context close to 30% of resource papers translate existing benchmarks to create language specific benchmarks Alabi et al. (2025). However, to better reflect native cultural identities and move away from the Western-centric concepts that often persist in translated benchmarks, some researchers have begun developing fully localized benchmarks Yu et al. (2025). This work complements manual localization and cultural adaptation efforts by reducing the scale and resources required to acquire such data. Just as traditional data augmentation methods improve model robustness without the need for entirely new data creation, our proposed approach similarly complements translation by enabling the creation of large-scale, culturally adapted datasets with minimal overhead.
@@
Automatic Entity based Augmentation. In addition to manual localization efforts carried out by volunteers, Ye et al. (2024) introduce a novel data augmentation technique that leverages large language models (LLMs). Specifically, they apply this method to enhance performance in few-shot Named Entity Recognition (NER) tasks, demonstrating that LLM-driven augmentations can serve as a valuable complement to human-curated resources.
@@
Mathematics and Cultural Entities.  Karim et al. (2025); Tomar et al. (2025) examined the influence of cultural context on mathematical problems by analyzing the impact of culturally grounded entities such as personal names and food items. However, one critical dimension that remains underexplored is the role of language itself and the biases that can be propagated through translation. Since the significant number of multilingual training and evaluation datasets are created via translation from high-resource languages Alabi et al. (2025), they often fail to capture the linguistic and cultural nuances of the target languages. Evaluating model robustness in the presence of culturally specific entities is a critical challenge, especially as LLMs are increasingly deployed in real-world applications. Native language speakers naturally refer to familiar names, organizations, and currencies from their own cultural context, making it essential for models to handle such variations reliably.
@@ 3.1 Pipeline
As shown in stage 7 of Table 1, our quality control measures ensure no overlapping or inconsistent entity replacements, enforce a single currency type per problem to avoid conversion errors, verify that localized outputs match the length of their direct translations, and check similarity between localized and translated text using the difflib library.333docs.python.org/3/library/difflib.html Since all languages share the same entity-replaced text ($x_{\text{ent}}$), human verification only requires comparing entities in the replacement dictionary. This prevents unnecessary additions by LLMs, maintains prompt consistency, and keeps outputs close to the original structure.
@@ 3.2 Datasets Used
Evaluation Dataset. AfriMGSM Adelani et al. (2024) is a manually translated benchmark spanning 18 languages sourced from MGSM Shi et al. (2022b). We found that $\approx$86% of the test set, includes at least one important entity. Using the English and manually translated pairs, we created a localized evaluation dataset by replacing these entities with culturally appropriate local alternatives while keeping the rest of the problem content unchanged. The correctness of the localized dataset was verified through manual inspection by the authors. Since the approach returns the unlocalized item if it fails, no additional noise is introduced.
@@
To ensure translation quality, we filtered the outputs using SSA-COMMET Li et al. (2025), a sentence-level semantic similarity metric that scores translation quality on a scale from 0 to 1, with 1 indicating perfect translation. Based on the score distribution shown in the Appendix F, we retained only translations with a COMMET score above 0.65. From these, we selected the top 1,500 examples per language for automatic localization and use as training data in our after localizing each. While intermediate experiments leverage Gemini-1.5-pro to evaluate the pipeline and assess its effectiveness due to cost constraints and to avoid unwanted behaviors, such as prompt caching [^4] https://ai.google.dev/gemini-api/docs/caching?lang=python, the final dataset is generated using Gemini-2.5-pro to ensure the highest quality standards.
Figure 3: Direct Translation(AfriMGSM) vs. Auto Localization(Localized-AfriMGSM) Numeric Match performance. We observe performance differences between translated and localized benchmark indicating a lack of robustness in LLM mathematical ability for real-life culturally localized MWP variants.
Figure 4: Effect of Cultural Entities on English Benchmarks. We investigate whether replacing default English entities with culturally specific entities ($x_{ent}$; see Table 1) influences model performance. The results show that across models and languages, the inclusion of local entities consistently shifts evaluation outcomes, indicating that cultural grounding plays a measurable role in benchmark performance.
@@ 4 Experimental Setup
Evaluated LLMs. In this work, we evaluated both open-source and closed-source models commonly studied in existing multilingual mathematical research. Our selection spans a range of model sizes, including Aya-expanse-32b, Deepseek-r1-distill-llama-70b, Gemma-2-9b-it, GPT-4o-mini, LLaMA-3-70B-Instruct and Mistrial-nemo-instruct-2407 Dang et al. (2024), The finetuning experiments are done only on LLaMA-3-8B-Instruct, Gemma-2-9b-it because of compute constraints.
Evaluation Metrics. Most existing work on mathematical word problems focuses on evaluating reasoning capabilities, where intermediate calculation steps are used as part of the input or as supervision for general benchmarking. However, due to the lack of multilingual reasoning benchmarks and the challenges involved in extracting step-by-step reasoning across 18 different languages, this research focuses solely on evaluating final answers. Several metrics are commonly used in mathematical evaluation research, including exact match, numeric match. In this work, we select the metrics that best align with the core objectives of our study. Exact Match (EM) evaluates whether the predicted answer exactly matches the reference answer as a string. Numeric Match (NM) checks whether the predicted numerical value matches the ground truth, ignoring differences in formatting such as units or punctuation after converting them into floating point data and account for errors between them. Moreover, for ensuring robustness and minimal noise via prompt selection, we adopted three prompt variations from Adelani et al. (2024) for our experiments. We further customized them to return only the final output without intermediate steps.
@@ 5 Results and Analysis
(RQ1) Can LLM-driven pipelines be used to localize translated benchmarks? To address the high cost associated with using human translators for creating localized versions of datasets, we introduce an automatic pipeline that generates culturally adapted versions of mathematical word problems. While this pipeline is not intended to replace human annotators, it serves as a valuable tool for producing augmented datasets that help improve the robustness of LLMs to entity-based variations in problems.
In our first set of experiments (please see Figure 2), we used Gemini-1.5-pro to generate localized versions of the dataset using both our six-stage localization pipeline and direct prompting via the Gemini API. Once the two versions of the localized dataset were generated, three annotators independently labeled each instance as either a valid or invalid localization. Additionally we wanted to show this method does not require the latest/largest models, allowing us to achieve stellar performance even with older models like Gemini-1.5-pro.
@@
Languages LLaMA-3-8B-Instruct Gemma-2-9b-it
$x_{\text{trans}}$ $x_{\text{ent}}$ $\hat{x}_{\text{loc}}$ all data # samples $x_{\text{trans}}$ $x_{\text{ent}}$ $\hat{x}_{\text{loc}}$ all data # samples
Hausa -0.13 -0.27 -1.20 0.80 5k -1.73 0.93 1.07 0.40 10k
Swahili 0.40 1.33 -0.53 0.67 1k -2.00 1.60 1.20 1.07 5k
Ewe -0.67 0.13 0.27 0.80 25k 0.27 -0.53 0.27 1.20 25k
Twi 0.67 -0.27 0.53 -0.93 10k -0.27 0.13 -0.53 0.67 1k
Wolof -0.27 -0.13 0.93 0.93 1k -0.40 -0.27 0.27 0.67 5k
Lingala -0.13 -0.67 0.27 0.67 1k 0.40 0.67 2.00 1.47 10k
Luganda -0.13 0.13 0.00 0.40 1k 0.13 -1.59 1.07 0.27 5k
Oromo -0.40 -0.80 -0.53 0.67 10k -0.27 -0.40 0.80 -1.07 1k
Shona -1.73 -0.27 -0.13 0.93 1k 2.27 -0.27 1.20 1.60 1k
Xhosa -0.27 -1.20 0.27 0.93 1k 0.67 0.00 1.07 1.87 1k
Yoruba -1.07 -0.67 -0.40 0.27 1k 0.00 0.13 1.20 0.13 1k
Kinyarwanda 0.27 -0.27 -0.40 1.73 1k 0.00 -1.60 0.67 0.67 25k
Zulu 0.27 -1.07 0.80 -0.80 1k -0.67 1.20 0.40 0.13 1k
Sotho -0.40 0.00 0.40 0.53 1k 0.00 -1.60 0.67 0.67 10k
Igbo 0.00 -0.27 0.13 0.67 1k -0.53 1.87 0.67 0.27 1k
Amharic 0.13 -1.20 1.07 0.93 1k -1.73 0.93 1.07 0.40 25k
French 1.07 0.93 0.67 -0.93 5k -1.07 -1.33 -0.93 0.53 10k
# Native Robust Lang. 6 4 10 14 5 8 15 16
Table 3: Native Robustness (Numeric Match $\Delta$). We report $\Delta_{\mathrm{NM}}=\mathrm{NM}_{\text{localized}}-\mathrm{NM}_{\text{translated}}$ across sampled data fine-tunings for translated data ($x_{\text{trans}}$), English entity–replaced data ($x_{\text{ent}}$), auto-localized data ($\hat{x}_{\text{loc}}$), and all data combined. Positive values indicate higher robustness on localized benchmarks, Negative values indicate stronger performance on English-centric benchmarks, and Yellow denotes no change. Languages are arranged in increasing resource order based on the FineWeb-2 dataset Penedo et al. (2025).
(RQ2) Does introducing local entities in translated benchmarks reveal performance disparities for MWPs in LRLs and English? We first analyze LRLs and then English language MWPs. Given that 30% of newly created African datasets are based on translated content Alabi et al. (2025), it is important to assess whether LLMs are overfitting to English-centric entities commonly preserved during translation. Such overfitting may lead to performance degradation when these entities are replaced with their native counterparts. In this work, we hypothesize that models should perform better when evaluated on data containing english entities because english centric bias that is found in most training and evaluation dataset.
Figure 3 illustrates the performance differences between models evaluated on translated benchmarks and those evaluated on automatically localized benchmarks, highlighting the impact of entity localization on model robustness. We can observe that the all models except Gemma-2-27b-it perform well on translated benchmarks. This shows the translated benchmarks tend to mislead the performance of models and the evaluation doesn’t directly simulate the real life usage when people use their local entities in the problems.
Gemma-2-27b-it demonstrates more consistent results across both native and English-centric benchmarks. At the language-specific level, we observe that French and Swahili, relatively higher-resource languages, show pronounced effects in the Deepseek-r1-distill-llama-70b and Mistrial-nemo-instruct-2407 models.
@@
Both LLaMA-3-70B-Instruct and GPT-4o-mini achieve higher accuracy on the localized (native-entity) variant than on the English benchmark, whereas the Aya-expanse-32b model’s performance lags behind. Gemma-2-27b-it generally tracks the localized variant more closely, despite occasional dips and overshoots. We have relatively similar trends across languages unlike LRL related experiments.
@@
Figure 3 presents the performance of models on different flavors of MWP data. From the multilingual dataset we created through translation ($x_{\text{trans}}$), English entity–replacement ($x_{\text{ent}}$), auto-localization ($\hat{x}_{\text{loc}}$), and a combination of all datasets, we randomly sampled subsets of 1k, 5k, 10k, and the full 25k examples for model training. We opted for various training set sizes since different languages require different volumes of training data to enhance model robustness across localized MWP versions.
Evaluation was then conducted on both the original translated datasets and our localized versions to assess performance differences achieved. This comparison helps us understand whether the inclusion of native cultural entities affects model behavior. We observe that both LLaMA-3-8B-Instruct and Gemma-2-9b-it models exhibit improved robustness to entity changes in several languages. For both models, the best performance is achieved when localization is combined with additional noisy data, indicating that diverse training sources can enhance generalization.
Looking at the translated ($x_{\text{trans}}$) and English entity–replaced ($x_{\text{ent}}$) data, Gemma-2-9b-it demonstrates stronger performance when local entities are present in the questions, whereas LLaMA-3-8B-Instruct exhibits a performance drop. Incorporating localized datasets in addition to language changes in the training set leads to improvements over purely translated benchmarks in both models, though Gemma-2-9b-it benefits more from this effect. Finally, combining all datasets yields models that are more robust to these variations.